OpenStudy (anonymous):

(2√7)² simplifed with work

OpenStudy (anonymous):

\[\sqrt{(2\sqrt{7)^2}} --> 2\sqrt{7}\]

OpenStudy (anonymous):

wait that doesn't make sense

OpenStudy (anonymous):

(2√7)² =2^2*7=4*7=28

OpenStudy (anonymous):

how do you get 4 times 7?

OpenStudy (anonymous):

2^2 =4 \[(\sqrt{7})^2=7\]